OpenStudy (anonymous):

a train travels some distance with a speed of 30 km/h. and then it returns to 45 km/h. calculate the average speed

OpenStudy (anonymous):

@ash2326 @AccessDenied @annas @Rohangrr

OpenStudy (anonymous):

the average speed is (s1*t1 + s2*t2)/(t1 + t2), where s1 and s2 are the speeds and t1 and t2 are the respective times.

OpenStudy (anonymous):

solve it further

OpenStudy (anonymous):

@lgbasallote @lalaly

OpenStudy (anonymous):

Let's say the distance is x. Total distance will be 2x. Time for first trip = x/30 Time for second trip = x/45 Average speed = 2x/(x/30 + x/45) = 2x/(x/18) = 36 km/h

OpenStudy (anonymous):

okay thats final solution then

OpenStudy (anonymous):

Why do you even need to ask if that is the final solution. Do it yourself and if you get that then it's right

OpenStudy (anonymous):

be quiet @Romero

OpenStudy (anonymous):

and I actually don't think that's the right answer btw

OpenStudy (anonymous):

Distance=D Total Distance =2D Time one way=D/V1=D/30 Time the other way: D/V2=D/45 Total Time =T1+T2=D/30+D/45 Average Speed=Total Distance/Total Time = 2D/(D/30+D/45)=2/(1/30+1/45)=36Km/h

OpenStudy (anonymous):

thank u all guys